# Graph Transformers: A Survey

### v1

@@ V Application Perspectives of Graph Transformers
Graph transformers are finding applications in various domains that involve interconnected data. This section delves into their applications for graph-related tasks, categorized by the level of analysis: node-level, edge-level and graph-level. Beyond these core tasks, graph transformers are also making strides in applications that handle text, images, and videos, where data can be effectively represented as graphs for analysis.
@@ V-A1 Protein Structure Prediction
In the field of bioinformatics, graph transformers have demonstrated substantial potential in Protein Structure Prediction (PSP) [154]. Gu et al. [155] introduced HEAL, which employs hierarchical graph transformers on super-nodes that imitate functional motifs to interact with nodes in the protein graph, effectively capturing structural semantics. Pepe et al. [156] used Geometric Algebra (GA) modelling to introduce a new metric based on the relative orientations of amino acid residues, serving as an additional input feature to a graph transformer, assisting in the prediction of the 3D coordinates of a protein. Chen et al. [157] proposed gated-graph transformers integrating node and edge gates within a graph transformer framework to regulate information flow during graph message passing, proving beneficial in predicting the quality of 3D protein complex structures. Despite the encouraging outcomes, various challenges persist, such as the complexity of protein structures, scarcity of high-quality training data, and substantial computational resource [158]. Further investigation is necessary to address these challenges and enhance the precision of these models.
@@ V-A2 Entity Resolution
Entity Resolution (ER) is a crucial task in data management that aims to identify and link disparate representations of real-world entities from diverse sources [159]. Recent research has highlighted the efficacy of graph transformers in ER. For example, Yao et al. [160] proposed Hierarchical Graph Attention Networks (HierGAT) integrating the self-attention mechanism and graph attention network mechanism to capture and leverage the relationships between different ER decisions, leading to substantial enhancements over conventional approaches. Ying et al. [161] extended the standard transformer architecture and introduced several straightforward yet powerful structural encoding techniques to enhance the modeling of graph-structured data. Despite facing challenges related to data complexity and structural information encoding, these techniques have exhibited promising outcomes in terms of enhanced performance, scalability, and accuracy. Furthermore, Dou et al.[162] proposed the Hybrid Matching Knowledge for Entity Matching (GTA) method improves the transformer for representing relational data by integrating additional hybrid matching knowledge acquired through graph contrastive learning on a specially designed hybrid matching graph. This approach has also demonstrated promising results by effectively boosting the transformer for representing relational data and surpassing existing entity matching frameworks.
@@ V-A3 Anomaly Detection
Graph transformers are valuable tools for anomaly detection, especially in dynamic graphs and time series data [65, 163]. They tackle key challenges like encoding information for unattributed nodes and extracting discriminative knowledge from spatial-temporal dynamic graphs. Liu et al [164] proposed TADDY, a transformer-based Anomaly Detection framework, enhancing node encoding to represent each node’s structural and temporal roles in evolving graph streams. Similarly, Xu et al. [165] proposed the Anomaly Transformer which uses an Anomaly-Attention mechanism to measure association discrepancy and employs a minimax strategy to enhance normal-abnormal differentiation. Chen et al. [166] proposed the GTA framework for multivariate time series anomaly detection incorporates graph structure learning, graph convolution, and temporal dependency modeling with a transformer-based architecture. Tuli et al. [167] developed TranAD, a deep transformer network for anomaly detection in multivariate time series data, showing efficient anomaly detection and diagnosis in modern industrial applications. Despite their effectiveness, further research is needed to enhance their performance and applicability across different domains.
@@ V-B1 Drug-Drug Interaction Prediction
Graph transformers have been increasingly employed in the prediction of Drug-Drug Interactions (DDIs) owing to their capability to adeptly model the intricate relationships between drugs and targets [170]. Wang et al. [109] proposed a method which uses a line graph with drug-protein pairs as vertices and a graph transformer network (DTI-GTN) for the purpose of forecasting drug-target interactions. Djeddi et al. [171] proposed a novel approach named DTIOG for the prediction of DTIs, leveraging a Knowledge Graph Embedding (KGE) strategy and integrating contextual information derived from protein sequences. Despite the encouraging outcomes, these approaches encounter challenges such as overlooking certain facets of the intermolecular information and identifying potential interactions for newly discovered drugs [172]. Nevertheless, findings from multiple studies indicate that graph transformers can proficiently anticipate DDIs and surpass the performance of existing algorithms [173].
@@ V-B2 Knowledge Graph Completion
In the domain of Knowledge Graph (KG) completion, the utilization of graph transformers has been extensively investigated [174]. Chen et al. [175] proposed a novel inductive KG representation model, known as iHT, for KG completion through large-scale pre-training. This model comprises an entity encoder and a neighbor-aware relational scoring function, both parameterized by transformers. The application of this approach has led to remarkable advancements in performance, with a relative enhancement of more than 25% in mean reciprocal rank compared to previous models. Liu et al. [176] introduced a generative transformer with knowledge-guided decoding for academic KG completion, which incorporates pertinent knowledge from the training corpus to provide guidance. Chen et al. [177] developed a hybrid transformer with multi-level fusion to tackle challenges in multimodal KG completion tasks. This model integrates visual and textual representations through coarse-grained prefix-guided interaction and fine-grained correlation-aware fusion modules.
V-B3 Recommender Systems
Graph transformers have been effectively utilized in recommender systems by combining generative self-supervised learning with a graph transformer architecture [178]. Xia et al. [179] used the generative self-supervised learning method to extract representations from the data in an unsupervised manner and utilized graph transformer architecture to capture intricate relationships between users and items in the recommendation system. Li et al. [180] introduced a new method for recommender systems that leverages graph transformers (GFormer). Their approach automates the self-supervision augmentation process through a technique called rationale-aware generative self-supervised learning. This technique identifies informative patterns in user-item interactions. The proposed recommender system utilizes a special type of collaborative rationale discovery to selectively augment the self-supervision while preserving the overall relationships between users and items. The rationale-aware self-supervised learning in the graph transformer enables graph collaborative filtering. While challenges remain in areas like graph construction, network design, model optimization, computation efficiency, and handling diverse user behaviors, experiments show that the approach consistently outperforms baseline models on various datasets.
@@ V-C1 Molecular Property Prediction
Graph transformers are powerful tools for molecular property prediction, utilizing the graph structure of molecules to capture essential structural and semantic information [181]. Chen et al. [182] proposed Algebraic Graph-Assisted Bidirectional Transformer (AGBT) framework which integrates complementary 3D molecular information into graph invariants, rectifying the oversight of three-dimensional stereochemical information in certain machine learning models. Li et al [144] utilized Knowledge-Guided Pre-training of Graph Transformer (KPGT) in a self-supervised learning framework which emphasizes the importance of chemical bonds and models the structural information of molecular graphs. Buterez et al. [183] proposed transfer learning with graph transformer to enhance molecular property prediction on sparse and costly high-fidelity data.
@@ V-C2 Graph Clustering
Graph transformers have been increasingly utilized in the field of Graph Clustering, offering innovative methodologies and overcoming significant challenges [184]. Yun et al. [185] proposed a graph transformer network to generate new graph structures for identifying useful connections between unconnected nodes on the original graph while learning effective node representation on the new graphs in an end-to-end fashion. Gao et al. [186] proposed a patch graph transformer (PatchGT) that segments a graph into patches based on spectral clustering without any trainable parameters and allows the model to first use GNN layers to learn patch-level representations and then use transformer to obtain graph-level representations. These methodologies have addressed issues such as the limitations of the local attention mechanism and difficulties in learning high-level information, leading to enhanced graph representation, improved model performance, and effective node representation.
@@ V-C3 Graph Synthesis
Graph transformers have been applied in the field of graph synthesis to improve graph data mining and representation learning. Existing graph transformers with Positional Encodings have limitations in node classification tasks on complex graphs, as they do not fully capture the local node properties [16]. To address this, Ma et al. [187] introduced the Adaptive Graph Transformer (AGT). This model tackles the challenge of extracting structural patterns from graphs in a way that is both effective and efficient. AGT achieves this by learning from two different graph perspectives: centrality and subgraph views. This approach has been shown to achieve state-of-the-art performance on real-world web graphs and synthetic graphs characterized by heterophily and noise. Jiang et al. [188] proposed an Anchor Graph Transformer (AGFormer) that leverages an anchor graph model to perform more efficient and robust node-to-node message passing for overcoming the computational cost and sensitivity to graph noises of regular graph transformers. Zhu et al. [122] developed a Hierarchical Scalable Graph Transformer (HSGT) which scales the transformer architecture to node representation learning tasks on large-scale graphs by utilizing graph hierarchies and sampling-based training methods.
@@ V-D1 Text Summarization
Text summarization is a crucial aspect of NLP which has been significantly advanced with the introduction of Graph transformers [190]. These models utilize extractive, abstractive, and hybrid methods for summarization. Extractive summarization selects and extracts key sentences or phrases from the original text to create a summary [191]. In contrast, abstractive summarization interprets the core concepts in the text and produces a concise summary. Hybrid summarization combines the advantages of both approaches [192]. Despite the progress, challenges remain in text comprehension, main idea identification, and coherent summary generation. Nevertheless, the application of graph transformers in text summarization has demonstrated promising outcomes in terms of summary quality and efficiency [193].
@@ V-D2 Image Captioning
Graph transformers have emerged as a potent tool within the domain of image captioning, offering a structured representation of images and efficiently processing them to produce descriptive captions [194]. Techniques such as Transforming Scene Graphs (TSG) leverage multi-head attention to architect graph neural networks for embedding scene graphs, which encapsulate a myriad of specific knowledge to facilitate the generation of words across various parts of speech [195]. Despite encountering challenges, such as training complexity, absence of contextual information, and lack of fine-grained details in the extracted features, graph transformers have exhibited promising outcomes. They have enhanced the quality of generated sentences and attained state-of-the-art performance in image captioning endeavors [196].
@@ V-D3 Image Generation
Graph transformers have been effectively utilized in image generation, as demonstrated by various research studies [197]. Sortino et al. [33] proposed a transformer-based method conditioned by scene graphs for image generation, employing a decoder to sequentially compose images. Zhang et al. [198] proposed StyleSwin which uses transformers in constructing a generative adversarial network for high-resolution image creation. Despite challenges like redundant interactions and the requirement for intricate architectures, these studies have exhibited promising outcomes in terms of image quality and variety [199, 200].
@@ V-D4 Video Generation
Graph transformers have been extensively applied in the field of Video Generation. Xiao et al. [201] proposed a Video Graph Transformer (VGT) model, which utilizes a dynamic graph transformer module to encode videos, capturing visual objects, their relationships, and dynamics. It incorporates disentangled video and text transformers for comparing relevance between the video and text. Wu et al. [202] proposed The Object-Centric Video Transformer (OCVT) which adopts an object-centric strategy to break down scenes into tokens suitable for a generative video transformer and understanding the intricate spatiotemporal dynamics of multiple interacting objects within a scene. Yan et al. [203] developed VideoGPT, which learns downsampled discrete latent representations of a raw video through 3D convolutions and axial self-attention. A GPT-like architecture is then used to model the discrete latent in a spatiotemporal manner using position encodings. Tulyakov et al. [204] proposed the MoCoGAN model which creates a video by mapping a series of random vectors to a sequence of video frames. Despite the challenges in capturing complex spatio-temporal dynamics in videos, these methodologies have exhibited promising outcomes across various facets of video generation, ranging from question answering to video summarization and beyond.
VI Open Issues and Future Directions
@@
VI-A Scalability and Efficiency
The scalability and efficiency of graph transformers pose considerable challenges due to their substantial memory and computational requirements especially when employing global attention mechanisms to deal with large-scale graphs [205, 114]. These challenges are further amplified in deep architectures which are susceptible to overfitting and over-smoothing. To address these issues, several potential strategies can be proposed:
@@ VI-A Scalability and Efficiency
    Developing efficient attention mechanisms, such as linear, sparse and low-rank attention, to reduce the complexity and memory usage of graph transformers.
@@
    Applying graph sparsification or coarsening techniques to decrease the size and density of graphs while maintaining their crucial structural and semantic information.
@@
    Exploring graph distillation or compression methods to create compact and effective graph transformer models, which are suitable for deployment on resource-limited devices.
@@
    Investigating regularization or normalization techniques, such as dropout, graph diffusion, convolution, and graph spectral normalization, to prevent overfitting and over-smoothing in graph transformer models.
VI-B Generalization and Robustness
Graph transformers often face challenges when it comes to generalizing to graphs that they have not encountered before or that fall outside of their usual distribution. This is especially true for graphs that have different sizes, structures, features, and domains. Additionally, graph transformers can be vulnerable to adversarial attacks and noisy inputs, which can result in a decline in performance and the production of misleading results. In order to improve the generalization and robustness of graph transformers, the following strategies could be taken into account:

### v2

@@ V Application Perspectives of Graph Transformers
Graph transformers are increasingly utilized in domains involving interconnected data. This section explores their applications in graph-related tasks, categorized by analysis level: node-level, edge-level, and graph-level. Additionally, graph transformers are advancing in text, image, and video applications, where data is represented as graphs for enhanced analysis.
@@ V-A1 Protein Structure Prediction
In bioinformatics, graph transformers have shown significant potential in Protein Structure Prediction (PSP) [142]. Gu et al. [143] introduced HEAL, which utilizes hierarchical graph transformers on super-nodes that mimic functional motifs to interact with nodes in the protein graph, capturing structural semantics. Pepe et al. [144] employed Geometric Algebra (GA) modeling to introduce a metric based on the relative orientations of amino acid residues, serving as an input feature to a graph transformer for predicting protein 3D coordinates. Chen et al. [145] proposed gated-graph transformers that integrate node and edge gates to regulate information flow during graph message-passing, aiding in predicting the quality of 3D protein complex structures. Despite promising results, challenges such as protein structure complexity, limited high-quality training data, and high computational demands persist [146]. Further research is needed to address these challenges and improve model accuracy.
@@ V-A2 Entity Resolution
Entity Resolution (ER) is essential in data management, aiming to identify and link disparate representations of real-world entities from various sources [147]. Recent studies have shown the effectiveness of graph transformers in ER. Yao et al. [148] introduced Hierarchical Graph Attention Networks (HierGAT), integrating self-attention and graph attention mechanisms to leverage relationships between different ER decisions, significantly improving over traditional methods. Ying et al. [149] extended the transformer architecture with structural encoding techniques to better model graph-structured data, demonstrating enhanced performance, scalability, and accuracy. Dou et al. [150] proposed the Hybrid Matching Knowledge for Entity Matching (GTA) method, which improves transformers for relational data by integrating hybrid matching knowledge through graph contrastive learning. This approach has shown promising results, surpassing existing entity matching frameworks. Despite challenges related to data complexity and structural information encoding, these methods have proven effective.
@@ V-A3 Anomaly Detection
Graph transformers are crucial for anomaly detection, particularly in dynamic graphs and time series data [60, 151]. They address challenges such as encoding information for unattributed nodes and extracting discriminative knowledge from spatial-temporal dynamic graphs. Liu et al. [152] introduced TADDY, a transformer-based anomaly detection framework that enhances node encoding to represent structural and temporal roles in evolving graph streams. Xu et al. [153] proposed the Anomaly Transformer, which uses an Anomaly-Attention mechanism to measure association discrepancy and a minimax strategy to differentiate normal and abnormal data. Chen et al. [154] developed the GTA framework for multivariate time series anomaly detection, incorporating graph structure learning, graph convolution, and temporal dependency modeling with a transformer-based architecture. Tuli et al. [155] created TranAD, a deep transformer network for anomaly detection in multivariate time series, demonstrating efficient anomaly detection and diagnosis in industrial applications. Although effective, further research is required to enhance performance and applicability across different domains [156].
@@ V-B1 Drug-Drug Interaction Prediction
Graph transformers have been increasingly utilized for predicting Drug-Drug Interactions (DDIs) due to their ability to model complex drug-target relationships [159]. Wang et al. [102] introduced a method using a line graph with drug-protein pairs as vertices and a graph transformer network (DTI-GTN) to predict drug-target interactions. Djeddi et al. [160] developed DTIOG, a novel approach leveraging Knowledge Graph Embedding (KGE) and contextual information from protein sequences to predict DTIs. Despite promising results, challenges remain, such as neglecting certain intermolecular information and identifying interactions for new drugs [161]. Nonetheless, multiple studies indicate that graph transformers can effectively predict DDIs, outperforming existing algorithms [162].
@@ V-B2 Knowledge Graph Completion
In Knowledge Graph (KG) completion, graph transformers have been extensively studied [163]. Chen et al. [164] proposed iHT, an inductive KG representation model for KG completion through large-scale pre-training, featuring an entity encoder and a neighbor-aware relational scoring function, both transformer-parameterized. This approach achieved over 25% improvement in mean reciprocal rank compared to previous models. Liu et al. [165] introduced a generative transformer with knowledge-guided decoding for academic KG completion, utilizing relevant knowledge from the training corpus. Chen et al. [166] developed a hybrid transformer with multi-level fusion for multimodal KG completion, integrating visual and textual representations via coarse-grained prefix-guided interaction and fine-grained correlation-aware fusion modules.
V-B3 Recommendation Systems
Graph transformers have been effectively employed in recommender systems by integrating generative self-supervised learning with a graph transformer architecture [167]. Xia et al. [168] utilized generative self-supervised learning to extract representations unsupervisedly and applied graph transformers to capture complex user-item relationships. Li et al. [169] introduced GFormer, which uses rationale-aware generative self-supervised learning to identify informative patterns in user-item interactions. This method automates self-supervision augmentation and preserves user-item relationships via collaborative rationale discovery. Despite challenges in graph construction, network design, model optimization, computation efficiency, and handling diverse user behaviors, experiments demonstrate that this approach consistently outperforms baseline models across various datasets [170].
@@ V-C1 Molecular Property Prediction
Graph transformers are highly effective for molecular property prediction, leveraging molecular graph structures to capture essential structural and semantic information [171]. Chen et al. [172] introduced the Algebraic Graph-Assisted Bidirectional Transformer (AGBT) framework, which integrates 3D molecular information into graph invariants, addressing the lack of stereochemical information in some models. Li et al. [132] utilized Knowledge-Guided Pre-training of Graph Transformers (KPGT) in a self-supervised learning framework, emphasizing chemical bonds and modeling molecular graph structures. Buterez et al. [173] proposed transfer learning with graph transformers to improve molecular property prediction on sparse, high-fidelity data.
@@ V-C2 Graph Clustering
Graph transformers are increasingly employed in Graph Clustering, offering innovative methods and addressing significant challenges [174]. Yun et al. [175] introduced a graph transformer network to generate new graph structures, identifying useful connections between previously unconnected nodes while learning effective node representations in an end-to-end manner. Gao et al. [176] proposed the Patch Graph Transformer (PatchGT), which segments a graph into patches via spectral clustering without trainable parameters, using GNN layers to learn patch-level representations and transformers for graph-level representations. These methodologies overcome limitations of local attention mechanisms and difficulties in learning high-level information, resulting in enhanced graph representation, improved model performance, and effective node representation.
@@ V-C3 Graph Synthesis
Graph transformers have advanced graph synthesis by enhancing graph data mining and representation learning. Traditional graph transformers with Positional encoding face limitations in node classification tasks on complex graphs due to inadequate capture of local node properties [16]. To address this, Ma et al. [177] introduced the Adaptive Graph Transformer (AGT), which effectively and efficiently extracts structural patterns from graphs by learning from centrality and subgraph views, achieving state-of-the-art performance on real-world web graphs and synthetic graphs with heterophily and noise. Jiang et al. [178] proposed the Anchor Graph Transformer (AGFormer), using an anchor graph model for efficient and robust node-to-node message-passing, overcoming computational costs and sensitivity to graph noise. Zhu et al. [114] developed the Hierarchical Scalable Graph Transformer (HSGT), scaling the transformer architecture to node representation learning tasks on large-scale graphs through graph hierarchies and sampling-based training methods.
@@ V-D1 Text Summarization
Text summarization, a crucial NLP task, has been significantly advanced by Graph transformers [180]. These models employ extractive, abstractive, and hybrid methods for summarization [181]. Extractive summarization selects key sentences or phrases from the original text [182], while abstractive summarization interprets core concepts to produce concise summaries. Hybrid summarization combines both approaches [183]. Despite challenges in text comprehension, main idea identification, and coherent summary generation, graph transformers have shown promising results in summary quality and efficiency [184].
@@ V-D2 Image Captioning
Graph transformers have become a powerful tool in image captioning, providing structured image representations and efficiently generating descriptive captions [185, 186]. Techniques like Transforming Scene Graphs (TSG) use multi-head attention to design graph neural networks that embed scene graphs, capturing extensive knowledge to aid word generation across various parts of speech [187]. Despite challenges like training complexity, lack of contextual information, and fine-grained details, graph transformers have shown promising results, improving sentence quality and achieving state-of-the-art performance in image captioning [188].
@@ V-D3 Image Generation
Graph transformers have been effectively utilized in image generation, as shown by various studies [189]. Sortino et al. [30] proposed a transformer-based method conditioned on scene graphs, employing a decoder for sequential image composition. Zhang et al. [190] introduced StyleSwin, using transformers in a generative adversarial network for high-resolution image creation. Despite challenges like redundant interactions and complex architectures, these studies have shown promising results in image quality and variety [120, 191].
@@ V-D4 Video Generation
Graph transformers have been extensively applied in video generation. Xiao et al. [192] introduced the Video Graph Transformer (VGT) model, using a dynamic graph transformer to encode videos by capturing visual objects, relationships, and dynamics, incorporating disentangled video and text transformers for relevance comparison. Wu et al. [193] proposed the Object-Centric Video Transformer (OCVT), employing an object-centric approach to tokenize scenes for a generative video transformer, understanding complex spatiotemporal dynamics. Yan et al. [194] developed VideoGPT, which uses 3D convolutions and axial self-attention to learn downsampled discrete latent representations of raw videos, modeling these latents with a GPT-like architecture. Tulyakov et al. [195] proposed MoCoGAN, mapping random vectors to video frame sequences. Despite challenges in capturing complex spatiotemporal dynamics, these methodologies have shown promising outcomes in various video generation tasks, from question answering to video summarization.
VI Design Guide for Graph Transformers
Developing effective graph transformers requires meticulous attention to graph-specific challenges and careful consideration of design choices. This guide outlines practical principles for designing graph transformers across various graph types and tasks.
  • Choose the appropriate type of graph transformer based on your graph data and tasks. Selecting a graph transformer requires a detailed understanding of the graph’s structure and scale. For ”simple, small” graphs, characterized by fewer than $10^{3}$ nodes and sparse connectivity (e.g., edge-to-node ratio <10<10), shallow transformers with fewer layers (e.g., 2–4 layers) may be sufficient to capture structural information efficiently. In contrast, ”complex, large” graphs, defined as having more than $10^{5}$ nodes, dense connectivity (edge-to-node ratio $\geq 10$), or rich attributes (e.g., temporal, weighted, or multi-relational edges), benefit from deeper models (e.g., 6–12 layers) capable of learning expressive representations. For dynamic or streaming graphs, where node or edge data evolves over time, scalable transformers leveraging techniques like subgraph sampling or incremental updates are essential [196]. Sparse or noisy graphs, such as those with missing edges or incomplete attributes, are effectively handled by pre-trained graph transformers, which leverage knowledge transfer and task-specific fine-tuning to improve robustness.
  • Design suitable structural and positional encodings for your graph data. Effective encoding of structural and positional information is crucial for transformers to capture graph-specific inductive biases. Techniques such as Laplacian eigenvector-based encodings, spectral embeddings, or random walk-based encodings (e.g., Node2Vec or DeepWalk embeddings) can encode global and local graph structure [197]. For weighted graphs, edge weights can be integrated into positional encodings using distance-weighted features or flow-based encodings. For directed graphs, direction-sensitive encodings such as signed Laplacian eigenvectors or edge direction indicators can capture asymmetry [198]. Incorporating these encodings into node or edge embeddings preserves structural informativeness and supports tasks such as graph classification and link prediction.
  • Optimize the self-attention mechanism for your graph data. The self-attention mechanism in graph transformers computes pairwise attention scores, which are computationally expensive for large graphs [199]. To address this, various optimization strategies can be employed:
    • Sampling: Node sampling (e.g., GraphSAGE [84]) or edge sampling (e.g., DropEdge [200]) reduces computation by operating on representative graph subsets while preserving global properties.
    • Sparsification: Techniques like top-$k$ sparsification retain only the most informative attention scores, reducing dense computations [201].
    • Partitioning: Graph partitioning methods (e.g., METIS [202]) divide the graph into smaller, manageable subgraphs processed independently, improving scalability [203].
    • Hashing: Methods such as locality-sensitive hashing approximate attention computations, enabling efficient processing of large node sets [204].
    • Regularization and normalization: Regularizers like Dropout and attention normalization techniques mitigate overfitting and stabilize training, addressing issues like oversmoothing and oversquashing in deep transformers.
    Each technique has trade-offs in terms of computational overhead, scalability, and performance fidelity, making the choice context-dependent.
  • Utilize pre-training techniques to enhance performance. Pre-training techniques leverage large-scale datasets or pre-trained models to improve task-specific performance. Methods such as graph contrastive learning (e.g., GraphCL [205]) enhance feature extraction by maximizing agreement between augmented graph views. Fine-tuning techniques adapt pre-trained models for domain-specific tasks, while distillation transfers knowledge from large, resource-intensive models to smaller, efficient ones. Pre-training is particularly beneficial for low-resource scenarios, where labeled data is scarce, enabling effective downstream task performance with limited additional training.
VII Open Issues and Future Directions
@@
VII-A Scalability and Efficiency
The scalability and efficiency of graph transformers face significant challenges due to high memory and computational demands, particularly with global attention mechanisms for large-scale graphs [206, 107]. These challenges are exacerbated in deep architectures prone to overfitting and over-smoothing. To mitigate these issues, several strategies are proposed:
@@ VII-A Scalability and Efficiency
    Developing efficient attention mechanisms, such as linear, sparse, and low-rank attention, to reduce complexity and memory usage.
@@
    Applying graph sparsification or coarsening techniques to decrease graph size and density while preserving key structural and semantic information.
@@
    Exploring graph distillation or compression methods to create compact and effective graph transformer models for deployment on resource-limited devices.
@@
    Investigating regularization or normalization techniques, such as dropout, graph diffusion, convolution, and graph spectral normalization, to prevent overfitting and over-smoothing.
VII-B Generalization and Robustness
Graph transformers often struggle to generalize to unseen or out-of-distribution graphs, especially those with different sizes, structures, features, and domains. They are also vulnerable to adversarial attacks and noisy inputs, leading to performance drops and misleading results. To enhance the generalization and robustness of graph transformers, consider the following strategies: